the sum of two numbers is 38. three times the smaller number exceeds the larger number by 30. find the numbers
Found 2 solutions by nerdybill, biancafranco86:
Answer by nerdybill(7384)

the sum of two numbers is 38. three times the smaller number exceeds the larger number by 30. find the numbers
.
Let x = smaller number
and y = larger number
.
from: "sum of two numbers is 38"
x+y = 38 (equation 1)
.
from: "three times the smaller number exceeds the larger number by 30" we get
3x = y+30 (equation 2)
.
Solve equation 1 for y:
x+y = 38
y = 38-x
Substitute above into equation 2:
3x = y+30
3x = 38-x+30
4x = 38+30
4x = 68
x = 17 (smaller number)
.
Larger number:
x+y = 38
17+y = 38
y = 21
.
The numbers are 17 and 21

Answer by biancafranco86(3)

Let x = smaller number
and y = larger number
.
from: "sum of two numbers is 38"
x+y = 38 (equation 1)
.
from: "three times the smaller number exceeds the larger number by 30" we get
3x = y+30 (equation 2)
.
Solve equation 1 for y:
x+y = 38
y = 38-x
Substitute above into equation 2:
3x = y+30
3x = 38-x+30
4x = 38+30
4x = 68
x = 17 (smaller number)
.
Larger number:
x+y = 38
17+y = 38
y = 21
.
The numbers are 17 and 21